# Controlling Large Electric Vehicle Charging Stations via User Behavior Modeling and Stochastic Programming

### v3

@@ I Introduction
${}^{\dagger}$ Work carried out while at Institut Polytechnique de Paris, Palaiseau, France
@@
The number of public electric vehicle (EV) charging stations (EVCS) has increased by 55% in 2022 [1], and this number is only expected to grow further as the EU and the US multiply incentives to support the development of publicly accessible fast-charging infrastructure [2, 3]. However, EV charging stations come with many challenges, both in their operation (e.g. load management, user access, payment, or compatibility) and in their integration into the existing electrical grids [4]. Additionally, EVCSs typically have to satisfy some power constraints, while satisfying the requests of their users and minimizing electricity costs. In this context, smart control strategies are required and will play a key role in scaling the EVCS network.
The difficulty of EVCS control comes from the many uncertainties involved in their operation. While users can be asked to indicate their intended parking time, they often, in practice, disconnect their vehicles much before; creating a need for session-end forecasts. As EVCSs frequently run on contracts having penalties for consumption peaks, planning the charge requires modeling the arrival time, session duration, and energy requests of future sessions. This turns the problem of EVCS control into a stochastic optimization problem, which can be tackled using different Stochastic Optimization approaches.
In [5], Kai-Wen Cheng et al. propose an EV charging algorithm focused on minimizing carbon emissions while satisfying user requests. However, they do not address the issue of early disconnections and instead assume that the exact disconnection times are provided as input by users and are respected. This contrasts with our approach, which relies on predicting arrival and departure time scenarios, allowing us to satisfy user requests even in case of early disconnection, and to better model the future EVCS load. Additionally, their algorithm lacks constraints on maximum power delivery, which we account for to respect real-life operating conditions [6, 7]. In [8], Tucker et al. do model a transformer capacity constraint and forecast EV departures. However, departure scenarios are determined deterministically at the connection time of each EV. Similarly, future EV arrival times are obtained from a pre-computed daily model obtained by aggregating past charging session data, independently of past slot activity. On the contrary, our approach models connection and departure times as a function of sojourn times and includes additional features such as charging slot indices to capture slot-specific behaviors. Furthermore, while their heuristic enforces a certain charging rate for the first connection hour, our algorithm relies on customer satisfaction modeling embedded in the objective function to ensure satisfaction even in case of early disconnection. The need for charging session forecasting is discussed in [9], where the necessity arises from correcting user-declared information about expected disconnection times and modeling future station load. The use of machine learning algorithms for predicting EV charging behavior — including session duration, start, and end times — has also been explored in [10] and [11], that advocate for the incorporation of external data such as weather and traffic alongside historical charging data. In our approach, we build upon these insights to advance the modeling of charging behavior by generating connection and disconnection scenarios for stochastic optimization. However, we maintain a focus on utilizing locally available data, which aligns more closely with our implementation and deployment constraints, similar to the ones discussed in [12].
@@
  • A modeling of the charging sessions’ end and start, formulated as a sojourn-dependent stochastic process, that improves the modeling approach presented in [12] and leverages user-provided information.
@@ II-C Charging dynamics
At each time step $t\in\mathcal{T}$, for each “active” slot $i$, considering a session with request $(k^{i}_{t_{0}},\delta^{i}_{t_{0}})$ starting at $t_{0}$, we also introduce a variable $z^{i}_{t}$ which stores the initial energy request of the session:
$z^{i}_{t}=k^{i}_{t_{0}}~{}\forall t,~{}t_{0}\leq t\leq t_{0}+\delta^{i}_{t_{0}}$ (2)
This variable will prove useful later to measure the satisfaction of the users.
Finally, we store the remaining number of time steps from the current time step $t$ before the announced end-of-parking time of slot $i$. We call this feature $m_{t}^{i}\in(\mathbb{N}\cup\{-\infty\})$. the variable takes value $\infty$- ∞ for charging sessions that are predicted using our methodology described in section IV-C.
@@ II-D Costs and objectives
Customer satisfaction. We define the satisfaction score $\chi^{i}_{t}$ of an “active” slot $i,~{}i<n$ using the ratio between the remaining energy $r_{i}^{t}$ to provide to the EV, and the associated initial energy requested $z^{i}_{t}=k^{i}_{t_{0}}$, where $t_{0}$ is the time step corresponding to the last connection:
@@
Electricity costs. The EVCS consumption at time $t\in\mathcal{T}$ is denoted by $c_{t}\in\mathbb{R}_{\geq 0}$ and corresponds to the sum of the consumption of its slots. The electricity cost at time step $t$ is denoted by $\zeta_{t}\in\mathbb{R}_{\geq 0}$ and is expressed as follows, with $p_{t}\in\mathbb{R}_{\geq 0}$ the electricity price at time $t$:
$\zeta_{t}$ $=p_{t}\times c_{t}$ (4)
$c_{t}$ $=\sum_{i\in S}e^{i}_{t}$ (5)
Grid import threshold overrun. At every time step $t\in\mathcal{T}$, a fixed penalty $\xi\in\mathbb{R}_{>0}$ is paid if the EVCS electricity consumption exceeds a threshold $\bar{c}\in\mathbb{R}_{>0}$, that is, if $c_{t}>\bar{c}$. At each time step $t$, the cost incurred by this threshold is then:
@@
where $\mathbf{1}_{>\bar{c}}:\mathbb{R}\to\{0,1\}$ is the indicator function that takes the value $1$ when its argument is above $\bar{c}$ and 00 otherwise.
@@ III Optimal control formulation
We formulate hereunder the problem as a generic Stochastic Optimal Control problem using standard notations used in the Control literature [14].
@@ III-A State, exogenous uncertainty and control variables
The state variable at time $t\in\mathcal{T}$ is denoted by $x_{t}\in\mathbb{X}=\big{(}\{0,1\}\times\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times(\mathbb{N}\cup\{-\infty\})\big{)}^{n}$ and is expressed as:
@@
The control (action) at time $t\in\mathcal{T}$ is denoted by $u_{t}=(e^{i}_{t})_{i=0}^{n-1}$ in the set $\mathbb{U}=(\mathbb{R}_{\geq 0})^{n}$ and gathers the power provided by each slot.
@@
The constraints and the dynamics are expressed as:
$g_{t}(x_{t},u_{t},w_{t})\leq 0$ (9)
$x_{t+1}=f_{t}(x_{t},u_{t},w_{t})$ (10)
where $f_{t}:(\mathbb{X},\mathbb{U},\mathbb{W})\to\mathbb{X}$ and $g_{t}:(\mathbb{X},\mathbb{U},\mathbb{W})\to\mathbb{R}^{n_{c}}$ denote respectively the dynamics function and the constraints, and where $n_{c}$ denotes the number of constraints.
We define the stage cost $L_{t}:(\mathbb{X},\mathbb{U})\to\mathbb{R}_{\geq 0}$ as a linear combination of the electricity cost (4), the grid threshold overrun penalty (6) and the satisfaction score (3), with weight given to the satisfaction denoted $\alpha\in\mathbb{R}_{>0}$. This weight gives a monetary value to the satisfaction of the users.
@@ III-B Controller definition and Multi-Stage Stochastic Programming policies
Our goal is to define a control strategy or policy. A policy is a collection $\Phi=\{\phi_{t_{0}}\}_{{t_{0}}\in\mathcal{T}}$ of mappings (one for each ${t_{0}}\in\mathcal{T}$), that each takes a current state $x_{t_{0}}\in\mathbb{X}$, and an uncertainty history $h_{t_{0}}=\{w_{t}\}_{t\leq{t_{0}}}$, and that returns a control action $u_{t_{0}}\in\mathbb{U}$ satisfying the problem constraints:
$\phi_{t_{0}}:(x_{t_{0}},h_{t_{0}})\mapsto u_{t_{0}}\in\{u\in\mathbb{U},~{}g_{t}(x_{t_{0}},u_{t_{0}},w_{t_{0}})\leq 0\}\$
At time step $t_{0}$, given a state $x_{t_{0}}$ and a history $h_{t_{0}}$, a controller of policy $\Phi$ solves the following Multi-Stage Stochastic Programming problem and returns one of its solutions:
@@ IV-A1 Two-stage Stochastic Programming (2S)
Two-stage Stochastic Programming uses a collection of $K\in\mathbb{N^{*}}$ samples indexed by $0\leq k<K$ denoted ($w_{t_{0}},\tilde{w}^{k}_{t_{0}+1},...,\tilde{w}^{k}_{t_{0}+R})_{k=0}^{K-1}$ drawn from the uncertainty distribution $\mathbb{P}(\mathbf{W}_{t_{0}},\ldots,\mathbf{W}_{t_{0}+R}|\mathbf{H}_{t_{0}}=h_{t_{0}})$, where the first variable of the sampled tuples, $w_{t_{0}}$, is observable and given by the history $h_{t_{0}}$ and is thus fixed. These samples are then assigned to $K^{\prime}\leq K$ clusters. The probability $\pi_{k}\in(0,1]$ of a cluster is then computed as the ratio between the number of samples associated with this cluster, and the total number of samples.
For each of the $K^{\prime}$ cluster centers, we find their closest sample, and we denote by ($w_{t_{0}},\hat{w}^{k}_{t_{0}+1},...,\hat{w}^{k}_{t_{0}+R})_{k=0}^{K^{\prime}-1}$ this new collection, that we then use to approximate the previously formulated problem by the following smaller and tractable optimization problem:
$\operatorname*{\text{arg}\,min}_{u_{t_{0}}\in U}$ $\min_{{(u^{k}_{t})_{\begin{subarray}{c}t_{0}<t\leq t_{0}+R\\ k<K^{\prime}\end{subarray}}}}\sum_{\begin{subarray}{c}t_{0}\leq t\leq t_{0}+R\\ k<K^{\prime}\end{subarray}}\pi_{k}\cdot L_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t})$ (13e)
s.t. $x^{k}_{t+1}=f_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t}),~{}\forall k<K^{\prime},~{}\forall t,~{}t_{0}\leq t<t_{0}+R$ (13f)
$g_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t})\leq 0,~{}\forall k<K^{\prime},~{}\forall t,~{}t_{0}\leq t\leq t_{0}+R$ (13g)
$x^{k}_{t_{0}}=x_{t_{0}},\quad\forall k<K^{\prime}$ (13h)
@@ IV-A2 Model Predictive Control (MPC)
$\operatorname*{\text{arg}\,min}_{u_{t_{0}}\in U}$ $\min_{{(u_{t})}_{t_{0}<t\leq t_{0}+R}}\sum_{t_{0}\leq t\leq t_{0}+R}L_{t}(x_{t},u_{t},\hat{w}_{t})$ (14a)
s.t. $x_{t+1}=f_{t}(x_{t},u_{t},\hat{w}_{t}),~{}\forall t,~{}t_{0}\leq t<t_{0}+R$ (14b)
$g_{t}(x_{t},u_{t},\hat{w}_{t})\leq 0,~{}\forall t,~{}t_{0}\leq t\leq t_{0}+R$ (14c)
@@ IV-B Uncertainty distribution modeling
In the following subsection, we explain how we model the uncertainty distribution, that we use to obtain the uncertainty samples used with 2S, and the uncertainty forecasts used with MPC.
@@ IV-B1 Session start and end
We model the charging session state transition of each slot $i,~{}i<n$ by a sojourn-dependent stochastic process $\mathbf{O}^{i}$ with two possible states, $\mathbf{O}^{i}(t)=0$ (inactive) or $\mathbf{O}^{i}(t)=1$ (active). This formulation is inspired by the Sojourn-time-dependent semi-markov switching processes introduced in [16].
Sojourn-time-dependent semi-markov switching processes are semi-markov processes in which the switching (i.e. transition between two distinct states) probability depends on the sojourn time (i.e. the time elapsed since the last switch). In our approach, we additionally condition the switching probability on the following set of features, which we directly derive from the uncertainty history $h_{t}$:
  • Sojourn time (Number of time steps since the last state transition). We keep track of the number of time steps since the last transition from “inactive” to “active” states, or from “inactive” to “active” states, in the variable $g_{t}^{i}\in\mathbb{N}$.
@@
The transition probability $p_{y\to z}$ from state $y\in\{0,1\}$ to state $z\in\{0,1\}$ of slot $i$ is thus defined as:
$p^{i}_{y\to z}=\mathbb{P}\bigg{(}\mathbf{O}(t)=z\bigg{|}\mathbf{O}(t-1)=y,\underbrace{g=g^{i}_{t-1}}_{\text{sojourn time}},l=l_{t-1},\tilde{i}=i\bigg{)}$ (15)
@@
We model the kWh request $k^{i}_{t}$ as a deterministic function denoted $\mathbf{k}$ of the sojourn time $g$, the time of the day $l$, and the slot index $i$: $\mathbf{k}:(\mathbb{N}\times\mathbb{N}\times\{0,1,...,n-1\})\to\mathbb{R}_{\geq 0}$.
@@ IV-C Uncertainty distribution approximation
  • Session starts. The first classifier $c_{1}$ is used to obtain the transition probabilities from “inactive” to “active” state, so as to obtain the session start uncertainty samples denoted $(a_{t_{0}},\tilde{a}_{t_{0}+1},...,\tilde{a}_{t_{0}+R})$. It approximates the distribution $p^{i}_{0\to 1}=\mathbb{P}(\mathbf{O}(t)=1|\mathbf{O}(t-1)=0,g=g^{i}_{t-1},l=l_{t-1},\tilde{i}=i)$ conditioned by the previous slot activation state, the sojourn time, the time and date, and the slot index.

### v4

@@ I Introduction
The number of public electric vehicle (EV) charging stations (EVCS) increased by 55% in 2022 [1], and this number is expected to grow as the EU and the US multiply incentives to support the development of publicly accessible fast-charging infrastructure [2, 3]. However, EV charging stations come with many challenges, both in their operation (e.g. load management, user access, payment, or compatibility) and in their integration into the existing electrical grids [4]. Additionally, EVCSs typically have to satisfy power constraints, while satisfying the requests of their users and minimizing electricity costs. In this context, smart control strategies are required and will play a key role in scaling the EVCS network.
The difficulty of EVCS control comes from the uncertainties involved in their operation. While users can be asked to indicate their intended parking time, they often, in practice, disconnect their vehicles much before; creating a need for session-end forecasts. As EVCSs frequently run on contracts having penalties for consumption peaks, planning the charge requires modeling the arrival time, session duration, and energy requests of future sessions. This turns the problem of EVCS control into a stochastic optimization problem, which can be tackled using different Stochastic Optimization approaches. In [5], Cheng et al. proposed an EV charging algorithm to minimize carbon emissions while meeting user demands. Unlike their approach, we predict arrival and departure times to handle early disconnections and better model EVCS load. We also consider constraints on maximum power delivery, addressing real-world conditions [6, 7]. Tucker et al. in [8] model transformer capacity and forecast EV departures but determine departure times deterministically. In contrast, our method models connection and departure times based on sojourn times, incorporating features like charging slot indices. We also ensure customer satisfaction through objective function embedding, even with early disconnections. The necessity for charging session forecasting is discussed in [9], advocating for machine learning-based prediction, including weather and traffic data. Our approach advances charging behavior modeling by generating connection and disconnection scenarios for stochastic optimization. More importantly, we do so by leveraging locally available data, while fully aligning with implementation constraints [10].
@@
  • A modeling of the charging sessions’ end and start, formulated as a sojourn-dependent stochastic process, which improves the modeling approach presented in [10] and leverages user-provided information.
@@ II-C Charging dynamics
At each time step $t\in\mathcal{T}$, for each “active” slot $i$, considering a session with request $(k^{i}_{t_{0}},\delta^{i}_{t_{0}})$ starting at $t_{0}$, we also introduce a variable $z^{i}_{t}$ which stores the initial energy request of the session, that is, $z^{i}_{t}=k^{i}_{t_{0}}~{}\forall t,~{}t_{0}\leq t\leq t_{0}+\delta^{i}_{t_{0}}$. This variable will prove useful later to measure the satisfaction of the users.
Finally, we store the remaining number of time steps from the current time step $t$ before the announced end-of-parking time of slot $i$. We call this feature $m_{t}^{i}\in(\mathbb{N}\cup\{-\infty\})$.
@@ II-D Costs and objectives
II-D1 Customer satisfaction
We define the satisfaction score $\chi^{i}_{t}$ of an “active” slot $i,~{}i<n$ at time $t$ using the ratio between the remaining energy $r_{i}^{t}$ to provide to the EV, and the associated initial energy requested $z^{i}_{t}=k^{i}_{t_{0}}$, where $t_{0}$ is the time step corresponding to the last connection:
@@ II-D1 Customer satisfaction
II-D2 Electricity costs
The EVCS consumption at time $t\in\mathcal{T}$ is denoted by $c_{t}\in\mathbb{R}_{\geq 0}$ and corresponds to the sum of the consumption of its slots. The electricity cost at time step $t$ is denoted by $\zeta_{t}\in\mathbb{R}_{\geq 0}$ and is expressed as follows, with $p_{t}\in\mathbb{R}_{\geq 0}$ denoting the electricity price at time $t$:
$\zeta_{t}=p_{t}\times c_{t},\qquad c_{t}=\sum_{i\in S}e^{i}_{t}$ (3)
II-D3 Grid import threshold overrun
At every time step $t\in\mathcal{T}$, a fixed penalty $\xi\in\mathbb{R}_{>0}$ is paid if the EVCS electricity consumption exceeds a threshold $\bar{c}\in\mathbb{R}_{>0}$, that is, if $c_{t}>\bar{c}$. At each time step $t$, the cost incurred by this threshold is expressed as follows, where $\mathbf{1}_{>\bar{c}}:\mathbb{R}\to\{0,1\}$ is the indicator function that takes the value $1$ when its argument is above $\bar{c}$ and 00 otherwise:
@@ III-A State, exogenous uncertainty and control variables
We formulate hereunder the problem as a generic Stochastic Optimal Control problem using standard notations used in the Control literature [12]. The state variable at time $t\in\mathcal{T}$ is denoted by $x_{t}\in\mathbb{X}=\big{(}\{0,1\}\times\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times(\mathbb{N}\cup\{-\infty\})\big{)}^{n}$:
@@
The control (action) at time $t\in\mathcal{T}$ is denoted by $u_{t}=(e^{i}_{t})_{i=0}^{n-1}$ in the set $\mathbb{U}=(\mathbb{R}_{\geq 0})^{n}$ and gathers the power provided by each slot. The dynamics and the constraints are expressed as $x_{t+1}=f_{t}(x_{t},u_{t},w_{t})$ and $g_{t}(x_{t},u_{t},w_{t})\leq 0$, where $f_{t}:(\mathbb{X},\mathbb{U},\mathbb{W})\to\mathbb{X}$ and $g_{t}:(\mathbb{X},\mathbb{U},\mathbb{W})\to\mathbb{R}^{n_{c}}$ denote respectively the dynamics function and the constraints, and where $n_{c}$ denotes the number of constraints. We define the stage cost $L_{t}:(\mathbb{X},\mathbb{U})\to\mathbb{R}_{\geq 0}$ as a linear combination of the electricity cost (3), the grid threshold overrun penalty (4) and the satisfaction score (2), with weight given to the satisfaction denoted $\alpha\in\mathbb{R}_{>0}$. This weight gives a monetary value to the satisfaction of the users:
@@ III-B Controller definition and Multi-Stage Stochastic Programming policies
Our goal is to define a control strategy or policy. A policy is a collection $\Phi=\{\phi_{t_{0}}\}_{{t_{0}}\in\mathcal{T}}$ of mappings (one for each ${t_{0}}\in\mathcal{T}$), that each takes a current state $x_{t_{0}}\in\mathbb{X}$, and an uncertainty history $h_{t_{0}}=\{w_{t}\}_{t\leq{t_{0}}}$, and that returns a control action $u_{t_{0}}$ solution to the following Multi-Stage Stochastic Programming problem:
@@ IV-A1 Two-stage Stochastic Programming (2S)
Two-stage Stochastic Programming uses a collection of $K\in\mathbb{N^{*}}$ samples indexed by $0\leq k<K$ denoted ($w_{t_{0}},\tilde{w}^{k}_{t_{0}+1},...,\tilde{w}^{k}_{t_{0}+R})_{k=0}^{K-1}$ drawn from the uncertainty distribution $\mathbb{P}(\mathbf{W}_{t_{0}},\ldots,\mathbf{W}_{t_{0}+R}|\mathbf{H}_{t_{0}}=h_{t_{0}})$, where the first variable of the sampled tuples, $w_{t_{0}}$, is observable and given by the history $h_{t_{0}}$ and is thus fixed. These samples are then assigned to $K^{\prime}\leq K$ clusters. The probability $\pi_{k}\in(0,1]$ of a cluster is then computed as the ratio between the number of samples associated with this cluster, and the total number of samples. For each of the $K^{\prime}$ cluster centers, we find their closest sample, and we denote by ($w_{t_{0}},\hat{w}^{k}_{t_{0}+1},...,\hat{w}^{k}_{t_{0}+R})_{k=0}^{K^{\prime}-1}$ this new collection, that we then use to approximate the previously formulated problem by the following smaller and tractable optimization problem:
$\operatorname*{\text{arg}\,min}_{u_{t_{0}}\in U}$ $\min_{{(u^{k}_{t})_{\begin{subarray}{c}t_{0}<t\leq t_{0}+R\\ k<K^{\prime}\end{subarray}}}}\sum_{\begin{subarray}{c}t_{0}\leq t\leq t_{0}+R\\ k<K^{\prime}\end{subarray}}\pi_{k}\cdot L_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t})$
s.t. $x^{k}_{t+1}=f_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t}),~{}\forall k<K^{\prime},~{}\forall t,~{}t_{0}\leq t<t_{0}+R$
$g_{t}(x^{k}_{t},u^{k}_{t},\hat{w}^{k}_{t})\leq 0,~{}\forall k<K^{\prime},\quad~{}\forall t,~{}t_{0}\leq t\leq t_{0}+R$
@@ IV-A2 Model Predictive Control (MPC)
$\operatorname*{\text{arg}\,min}_{u_{t_{0}}\in U}$ $\min_{{(u_{t})}_{t_{0}<t\leq t_{0}+R}}\sum_{t_{0}\leq t\leq t_{0}+R}L_{t}(x_{t},u_{t},\hat{w}_{t})$
s.t. $x_{t+1}=f_{t}(x_{t},u_{t},\hat{w}_{t}),~{}\forall t,~{}t_{0}\leq t<t_{0}+R$
$g_{t}(x_{t},u_{t},\hat{w}_{t})\leq 0,\quad~{}\forall t,~{}t_{0}\leq t\leq t_{0}+R$
@@ IV-B Uncertainty distribution modeling
In the following subsection, we explain the uncertainty distribution modeling necessary to obtain the uncertainty samples used with 2S, and the forecasts used with MPC.
@@ IV-B1 Session start and end
We model the charging session state transition of each slot $i,~{}i<n$ by a sojourn-dependent stochastic process $\mathbf{O}^{i}$ with two possible states, $\mathbf{O}^{i}(t)=0$ (inactive) or $\mathbf{O}^{i}(t)=1$ (active). This formulation is inspired by the Sojourn-time-dependent semi-markov switching processes introduced in [14]. Sojourn-time-dependent semi-markov switching processes are semi-markov processes in which the switching (i.e. transition between two distinct states) probability depends on the sojourn time (i.e. the time elapsed since the last switch). In our approach, we additionally condition the switching probability on the following set of features, which we directly derive from the uncertainty history $h_{t}$:
  • Sojourn time. We store the number of time steps since the last transition from “inactive” to “active” states, or from “inactive” to “active” states, in the variable $g_{t}^{i}\in\mathbb{N}$.
@@
The transition probability $p_{y\to z}$ from state $y\in\{0,1\}$ to state $z\in\{0,1\}$ of slot $i$ is thus defined as: $p^{i}_{y\to z}=\mathbb{P}\big{(}\mathbf{O}(t)=z\big{|}\mathbf{O}(t-1)=y,\underbrace{g=g^{i}_{t-1}}_{\text{sojourn time}},l=l_{t-1},\tilde{i}=i\big{)}$
@@
We model the kWh request $k^{i}_{t}$ as a deterministic function $\mathbf{k}$ of the sojourn time $g$, the time of the day $l$, and the slot index $i$: $\mathbf{k}:(\mathbb{N}\times\mathbb{N}\times\{0,1,...,n-1\})\to\mathbb{R}_{\geq 0}$.
@@ IV-C1 Session starts
The first classifier $c_{1}$ is used to obtain the transition probabilities from “inactive” to “active” state, so as to obtain the session start uncertainty samples denoted $(a_{t_{0}},\tilde{a}_{t_{0}+1},...,\tilde{a}_{t_{0}+R})$. It approximates the distribution $p^{i}_{0\to 1}=\mathbb{P}(\mathbf{O}(t)=1|\mathbf{O}(t-1)=0,g=g^{i}_{t-1},l=l_{t-1},\tilde{i}=i)$ conditioned by the previous slot activation state, the sojourn time, the time and date, and the slot index.
IV-C2 Kwh requests